OpenStudy (anonymous):

Which number should be placed in the blank to balance the equation 4Fe + ___ O2 → 2Fe2O3? 2 3 4 5

OpenStudy (anonymous):

2 *O3 = how many oxygen atoms?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

let me rephrase that how many oxygen atoms are in 3Fe2O3?

OpenStudy (anonymous):

... i actually don't know that's why I'm asking for help xP how do you find it? O.o

OpenStudy (anonymous):

http://openstudy.com/code-of-conduct Fe2O3 has 3 oxygen atoms in it so 2(Fe2O3) has how many oxygen atoms in it?

OpenStudy (anonymous):

I beleive its 5?

OpenStudy (anonymous):

2*3=?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

so how many oxygen atoms are on the right side?

OpenStudy (anonymous):

... there are 6 oxygen atoms on the right side........ 2Fe2O3

OpenStudy (anonymous):

wait but how do I find the answer then? lol sorry x3 I was going off of the answers it gave me :P

OpenStudy (anonymous):

\[2Fe_2O_3\] since we know there's a total of 6 oxygen atoms on the right sides this means there must be a total of 6 oxygen atoms on the left side \[AO_2\] so we need to determine the number of oxygen molecules since an oxygen molecules contains 2 oxygen atoms we can conclude A*2=6 A=?

OpenStudy (anonymous):

ooooooooooooooohhhhhhhhhhhhhhhh lmao xD A=3 :D

OpenStudy (anonymous):

Thanks sooooo much ;)

OpenStudy (anonymous):

this means there can be a total of 3 O2 in order to make the entire thing balanced